Write the slope-intercept form of the equation of the line that passes through the point (-2,4)and is perpendicular to the line with equation 2x-y+12=0.

Answer by mananth(16946)

y = 2 x + 12
Compare this equation with y=mx+b, m= slope & b= y intercept
slope m = 2

The slope of a line perpendicular to the above line will be the negative reciprocal -0.5
m1*m2=-1
The slope of the required line will be - 1/2

m= - 1/2 ,point ( -2 , 4 )
Find b by plugging the values of m & the point in
y=mx+b
4 = 1.00 + b
b= 3
m= - 1/2
The required equation isy =- 1/ 2x + 3
m.ananth@hotmail.ca